One type of antifreeze is 40% glycol, and another type of antifreeze is 60% glycol. How much of each kind should be used to make 100 gallons of antifreeze that is 48% glycol? I know answer is 60 gallons of 40% and 40 gallons of 60% but need to see the equation and steps to solve to get these answers? Thank you!
Answer by dkppathak(439)

One type of antifreeze is 40% glycol, and another type of antifreeze is 60% glycol. How much of each kind should be used to make 100 gallons of antifreeze that is 48% glycol?
let 40% glycole added is X gallon than pure glycole will be 40x/100=0.4X gallon
similarly let 60% glycol is Y gallon than pure glycol=60y/100=0.6Y gallon
pure glycol in 100 gallon should be 48% means pure glycol will be 48 gallon
as per given conditions
X+Y=100 (1)
0.4X+0.6Y =48(2) or 4x+6y=480 (2)
by multiplying (1) 4 and solving by elimination method
4X+4Y =400
4x+6y=480
Y=40 gallon
and X=60 gallon
Answer 60 gallon of 40% and 40 gallon of 60% glycol to be added for 48% glycol